ONLY 45 MINUTES LEFT!

Question 4 options:

LM.92 XYZ Company is rethinking the way it ships to its 58 customers in another city 300 miles away.

Current Shipping/Delivery Method

They currently hire an LTL (less-than-truckload) carrier to pick up and deliver these shipments. Each customer order shipped via LTL carrier costs $139.

Alternate Shipping/Delivery Method

A 3PL (third-party logistics provider) has approached XYZ Company and suggested that they make full truckload (TL) shipments from their facility to the 3PL's warehouse in the customers' city. The 3PL would then break the bulk shipment (TL or truckload shipment) into individual customer orders to be shipped locally by an LTL carrier. The relative data for this alternative shipping method are as follows:

  • Full TL shipment cost (300 miles) = $835
  • Average order weight = 565 lbs.
  • Warehouse break-bulk fee (per 100 lbs., a.k.a. per "hundred weight") = $13
  • Local LTL delivery fee = $43

What is the total cost of delivering to all customers via LTL carrier (current method)? (Display your answer as a whole number.)

What is the total cost of shipping to all customers using the alternate shipping/delivery method? (Display your answer as a whole number.)

At what cost of breaking bulk per hundred weight would the cost of these two methods be the same? (Display your answer to two decimal places.)
Answer #1

a)  total cost of delivering to all customers via LTL carrier = Total customers * Cost per customer = 58 * 139 = $8062

b)  total cost of shipping to all customers using the alternate shipping/delivery method = Full TL shipment cost +Local LTL delivery fee* Number of local deliveries (58) + Warehouse break-bulk fee * Average order weight/100 * Number of orders (58)

=835+43*58+13*565/100*58= $7589

c) At this cost, cost of both methods will be equal. Let the cost of bulk breaking be X.

$8062 = 835 +43*58+ X * 565/100*58

327.7 * X = 4733

X = $14.44

At this cost of breaking bulk per hundredweight, both methods have the same cost.

Then,